OpenStudy (anonymous):

58. A block of ice loses one third of its weight every hour. After 1 hour, the block weighs 60 pounds. How much did the block weigh to begin with? A. 90 lbs B. 120 lbs C. 100 lbs

OpenStudy (anonymous):

a is the answer

OpenStudy (anonymous):

figured it out :D

OpenStudy (jagatuba):

yes how did you figure it out? or did you guess?

OpenStudy (jagatuba):

nm

OpenStudy (anonymous):

no :D

OpenStudy (anonymous):

1/3 of 90 = 30 - 90 = 60

OpenStudy (jagatuba):

that is working back wards, but yes. What if you didn't have the choices to choose from?

OpenStudy (anonymous):

well then i dont know how i would have got it

OpenStudy (jagatuba):

Well since only one hour has passed we know that 2/3 of the ice is left right?

OpenStudy (jagatuba):

So we could ask a question that we can convert into an equation: Two-thirds of what is 60? OR 2/3 * X = 60 Make sense?

OpenStudy (anonymous):

yes :D

OpenStudy (jagatuba):

Now we just solve for X X = 60/(2/3)

OpenStudy (anonymous):

90

OpenStudy (jagatuba):

yes